### v2

@@ Abstract
Hyperspectral images (HSIs) are often contaminated by a mixture of noises such as Gaussian noise, dead lines, stripes, and so on. In this paper, we propose a novel approach for HSI denoising and destriping, called MLTL2p, which consists of an orthogonal constrained minimization model and an iterative algorithm with convergence guarantees. The model of the proposed MLTL2p approach is built based on a new sparsity-enhanced Multi-scale Low-rank Tensor regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. The multi-scale low-rank regularization for HSI denoising utilizes the global and local spectral correlation as well as the spatial nonlocal self-similarity priors of HSIs. The corresponding low-rank constraints are formulated based on independent higher-order singular value decomposition with sparsity enhancement on its core tensor to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for HSI destriping is extended from the matrix $\ell_{2,p}$ norm. A proximal block coordinate descent algorithm is proposed in the MLTL2p approach to solve the resulting nonconvex nonsmooth minimization with orthogonal constraints. We show any accumulation point of the sequence generated by the proposed algorithm converges to a first-order stationary point, which is defined using three equalities of substationarity, symmetry, and feasibility for orthogonal constraints. In the numerical experiments, we compare the proposed method with state-of-the-art methods including a deep learning based method, and test the methods on both simulated and real HSI datasets. Our proposed MLTL2p method demonstrates outperformance in terms of metrics such as mean peak signal-to-noise ratio as well as visual quality.
@@ 1 Introduction
HSIs are often contaminated by Gaussian noise, impulse noise, stripe noise, dead line noise, and so on. Mathematically, a noisy HSI $\mathcal{D}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ can be expressed as
@@
where $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the clean HSI, $\mathcal{S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the sparse noises such as impulse noise, stripe noise and dead line noise, and $\mathcal{N}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents Gaussian noise.
@@
To remove Gaussian noise in HSIs, many methods have been proposed. Conventional 2-D methods [13, 16], processing HSIs band by band, do not fully utilize the strong correlation between adjacent bands. 3-D methods such as Block-matching 4-D filtering (BM4D) [28], spectral-spatial adaptive hyperspectral total variation (SSAHTV) model [44], and sparse representation methods [49, 39] incorporate both spatial and spectral information and outperform conventional methods. However, those methods may fail to remove non-Gaussian noise.
@@
In real-world scenarios, HSIs are often contaminated by more than one type of noises due to atmospheric effects and instrument noise. Various methods have been proposed to remove the mixed noise, including low-rank matrix based methods, low-rank tensor based methods, and deep learning based methods. Low-rank matrix based methods [20, 46] reshape an HSI into a matrix and impose low-rankness on the reshaped HSI. Zhang et al. [47] formulated the HSI denoising problem as a low-rank matrix factorization problem and solved it by the “Go Decomposition” algorithm; Zhang et al. [46] proposed a double low-rank matrix decomposition which utilized the $\ell_{1}$ norm for the impulse noise and the matrix nuclear norm for stripes, and adopted augmented Lagrangian method (ALM) to solve the model; Yang et al. [42] also used double low-rankness, but added spatial-spectral total variation (SSTV) to the model and performed the decomposition on full band blocks (FBBs) rather than the entire HSI.
@@
Low-rank tensor based methods [22, 14, 30, 18, 50] view HSIs as tensors and perform tensor low-rank decompositions while preserving the spatial-spectral correlations. Wang et al. [36] used Tucker tensor decomposition and an anisotropic SSTV regularization to characterize the piece-wise smooth structures of the HSI; Chen et al.  [12] proposed a low-rank tensor decomposition (LRTD) method, which utilized the higher-order singular value decomposition (HOSVD) for low-rankness and the matrix $\ell_{2,1}$ norm for characterizing the stripes, and adopted ALM for solving the optimization model; Cao et al. [6] proposed a subspace-based nonlocal low-rank and sparse factorization (SNLRSF) method for removing mixed noise in HSI, which conducted nonlocal low-rank factorization via successive singular value decomposition (SVD); Xiong et al. [40] proposed the LRTFL0 method using low-rank block term decomposition and spectral-spatial $\ell_{0}$ gradient regularization to achieve gradient smoothness.
@@
For removing Gaussian noise, the exiting low-rank based methods mainly utilize global spectral correlation of HSIs, and do not use its nonlocal self-similarity prior [5], except for SNLRSF. This limits the performance of restoring clean HSIs. Also, many of those methods use ALM to solve their low-rank regularization based models without providing any convergence analysis. In their ALM, the augmented Lagrangian subproblem is solved by alternating minimization via one Gauss–Seidel-type iteration, instead of an exact joint minimization. As a result, the convergence is not guaranteed for these nonconvex nonsmooth optimization problems with three or more blocks [10].
@@
Recently, some deep neural networks [4, 45, 38] have been proposed to denoise HSIs. Quasi-recurrent neural networks (QRNN) [4] combined recurrent neural networks (RNN) with convolutional neural networks (CNN); the 3-D version of QRNN (QRNN3D) [38] can effectively embed the correlation of HSIs; CNN can also be used as a denoiser in a plug-and-play fashion for HSI denoising [34]. However, the performance of these methods may vary significantly depending on factors such as the datasets used, noise characteristics, and the size of the input data.
@@
(1) $\begin{split}\min_{\begin{subarray}{c}\mathcal{S},[{X}_{1}],[{X}_{2}],\\ [{X}_{3}],[\mathcal{G}],\mathcal{L}\end{subarray}}\quad&\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\|\mathcal{S}\|^{p}_{2,p}+\|[\mathcal{G}]\|_{1,w}\\ &+\frac{\delta}{2}\|\mathbf{R}(\mathcal{L})-[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]\|^{2}_{F}\\ \mbox{s.t. }&[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\end{split}$
@@
  • $\mathcal{S},\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $[\mathcal{G}]\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}\times N}$ and $[X_{i}]\in\mathbb{R}^{m_{i}\times n_{i}\times N}$, $m_{i}\geq n_{i}$, $i=1,2,3$;
@@
    (2) $\|\mathcal{S}\|_{2,p}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\|{s}_{:i_{2}i_{3}}\|^{p}_{2}\right)^{\frac{1}{p}}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\left(\sum_{i_{1}=1}^{I_{1}}{s}_{i_{1}i_{2}i_{3}}^{2}\right)^{\frac{p}{2}}\right)^{\frac{1}{p}},$
    and the tensor $\ell_{2,p}$ norm is exactly equal to the matrix $\ell_{2,p}$ norm of the unfolding matrix of ${\mathcal{S}}$ along the first dimension, i.e., $\|{\mathcal{S}}\|_{2,p}=\|{S}_{(1)}\|_{2,p}$;
  • $\|[\mathcal{G}]\|_{1,w}$ denotes the weighted tensor (component-wise) $\ell_{1}$ norm for a fourth order tensor $[\mathcal{G}]$ defined by
    (3) $\|[\mathcal{G}]\|_{1,w}=\sum_{j=1}^{N}w_{j}\|[\mathcal{G}]^{(j)}\|_{1},$
    with a weight vector $w\in\mathbb{R}^{N}_{++}$;
  • $\mathbf{R}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ denotes the block extraction operator that extracts blocks and forms a fourth order tensor, for example, the nonlocal similar FBB tensor, which can be regularized to exhibit low-rank properties;
  • $[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]$ denotes an independent 3-D HOSVD with $[\mathcal{G}]$ being the independent core tensors and $[X_{i}]$ being the independent mode-$i$ factor matrices, $i=1,2,3$, and $[X_{i}]$ is independently orthogonal, i.e., $[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}]$, with $\mathbf{I}_{n_{i}}$ representing the identity matrix of size $n_{i}\times n_{i}$.
Model (1) is a nonconvex nonsmooth minimization problem with orthogonal constraints. In particular, the first term of model (1) is a data fidelity term to remove Gaussian noise, the second term is a group sparsity measure to remove sparse noises with linear structures, and the last two terms are the sparsity-enhanced nonlocal low-rank tensor regularization terms.
@@
  • We propose a novel model to remove mixed noise in HSIs using a new sparsity-enhanced low-rank regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. For removing Gaussian noise, a tensor formed via the extraction operator $\mathbf{R}$ is regularized on its low-rankness using independent 3-D HOSVD with sparsity enhancement on its core tensors to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for removing the dead lines and stripes is extended from the matrix $\ell_{2,p}$ norm. Some new mathematical results on finding the proximal operator of the tensor $\ell_{2,p}$ norm are also presented. We show that model (1) has a nonempty and bounded solution set.
@@ 2.1 Notations
First, we introduce the notations for tensors. For a third order tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we let $x_{i_{1}i_{2}i_{3}}$ denote its $(i_{1},i_{2},i_{3})$-th entry, let $x_{:i_{2}i_{3}}$ denote its $(i_{2},i_{3})$-th mode-$1$ fiber and let $X_{::i_{3}}$ denote its $i_{3}$-th frontal slice. And $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$ means all its entries are positive. The mode-$k$ unfolding of a third order tensor $\mathcal{X}$ is denoted as $X_{(k)}={\rm unfold}_{(k)}(\mathcal{X})$, which is the process to linearize all indexes except index $k$. The dimensions of $X_{(k)}$ are $I_{k}\times\prod^{3}_{j=1,j\neq k}I_{j}$. An element $x_{i_{1}i_{2}i_{3}}$ of $\mathcal{X}$ corresponds to the position of $(i_{k},j)$ in matrix $X_{(k)}$, where $j=1+\sum^{3}_{l=1,l\neq k}(i_{l}-1)\prod^{l-1}_{m=1,m\neq k}I_{m}$. The inverse process of the mode-$k$ unfolding of a tensor $\mathcal{X}$ is denoted by $\mathcal{X}={\rm fold}_{(k)}(X_{(k)})$. An illustration of the unfoldings is presented in Fig. 2.
Figure 2: The unfoldings of tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$
@@
Next, we provide the definitions for (limiting) subdifferentials and proximal operators. Let $f:\mathbb{R}^{d}\to(-\infty,+\infty]$ be a proper and lower semicontinuous function with a finite lower bound. The (limiting) subdifferential of $f$ at $x\in\operatorname{dom}f:=\{x\in\mathbb{R}^{d}|f(x)<\infty\}$, denoted by $\partial f(x)$, is defined as
$\partial f(x)\!:=\{u\in\mathbb{R}^{d}|\exists x^{k}\to x,f(x^{k})\to f(x)\text{ and }u^{k}\to u\text{ with }u^{k}\in\hat{\partial}f(x^{k})\mbox{ as }k\to\infty\},$
where ^$\hat{\partial}f(x)$ denotes the Fréchet subdifferential of $f$ at $x\in\operatorname{dom}f$, which is the set of all $u\in\mathbb{R}^{d}$ satisfying
(4) $\begin{split}\underset{y\neq x,y\rightarrow x}{\lim\inf}\frac{f(y)-f(x)-\langle u,y-x\rangle}{\|y-x\|}\geq 0.\end{split}$
One can also observe that $\{u\in\mathbb{R}^{d}|\exists x^{k}\to x,f(x^{k})\to f(x)\text{ and }u^{k}\to u\text{ with }u^{k}\in\partial f(x^{k})\mbox{ as }k\to\infty\}\subseteq\partial f(x)$. Next, the proximal operator of $f$ with parameter $\lambda>0$ evaluated at $x\in\mathbb{R}^{d}$, denoted as $\text{prox}_{\lambda f}(x)$, is defined as
$\text{prox}_{\lambda f}(x):=\underset{u\in\mathbb{R}^{d}}{\arg\!\min}\left[f(u)+\frac{1}{2\lambda}\|u-x\|_{2}^{2}\right].$
Note that $\text{prox}_{\lambda f}$ is a single-valued map, if $f$ is a convex function. However, when $f$ is nonconvex, $\text{prox}_{\lambda f}(x)$ may have multiple points.
Also, we set $\mathbb{S}_{m,n}:=\{X\in\mathbb{R}^{m\times n}|X^{\top}X=\mathbf{I}_{n}\}$ as the Stiefel manifold with $m\geq n$ and set $\mathcal{T}_{X}\mathbb{S}_{m,n}:=\{Y\in\mathbb{R}^{m\times n}|Y^{\top}X+X^{\top}Y=0\}$ as the tangent space of Stiefel manifold at $X\in\mathbb{R}^{m\times n}$. We also set the Riemannian metric on Stiefel manifold as the metric induced from the Euclidean inner product. Then according to [1], the Riemannian gradient of a smooth function $f$ at $X$ is given by
$\operatorname{grad}f(X):=\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}(\nabla f(X)),$
where $\operatorname{Proj}_{\mathcal{T}_{X}\mathbb{S}_{m,n}}(Y):=(\mathbf{I}_{m}-XX^{\top})Y+\frac{1}{2}X(X^{\top}Y-Y^{\top}X)$.

### v3

@@ Abstract
Hyperspectral images (HSIs) are often contaminated by a mixture of noise such as Gaussian noise, dead lines, stripes, and so on. In this paper, we propose a multi-scale low-rank tensor regularized $\ell_{2,p}$ (MLTL2p) approach for HSI denoising and destriping, which consists of an orthogonal constrained minimization model and an iterative algorithm with convergence guarantees. The model of the proposed MLTL2p approach is built based on a new sparsity-enhanced Multi-scale Low-rank Tensor regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. The multi-scale low-rank regularization for HSI denoising utilizes the global and local spectral correlation as well as the spatial nonlocal self-similarity priors of HSIs. The corresponding low-rank constraints are formulated based on independent higher-order singular value decomposition with sparsity enhancement on its core tensor to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for HSI destriping is extended from the matrix $\ell_{2,p}$ norm. A proximal block coordinate descent algorithm is proposed in the MLTL2p approach to solve the resulting nonconvex nonsmooth minimization with orthogonal constraints. We show any accumulation point of the sequence generated by the proposed algorithm converges to a first-order stationary point, which is defined using three equalities of substationarity, symmetry, and feasibility for orthogonal constraints. In the numerical experiments, we compare the proposed method with state-of-the-art methods including a deep learning based method, and test the methods on both simulated and real HSI datasets. Our proposed MLTL2p method demonstrates outperformance in terms of metrics such as mean peak signal-to-noise ratio as well as visual quality.
@@ 1 Introduction
HSIs are often contaminated by additive white Gaussian noise, impulse noise, stripe noise, dead line noise, and so on. Mathematically, a noisy HSI $\mathcal{D}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ can be expressed as
@@
where $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the clean HSI, $\mathcal{S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the sparse noise such as impulse noise, stripe noise and dead line noise, and $\mathcal{N}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents additive white Gaussian noise.
@@
To remove additive white Gaussian noise in HSIs, many methods have been proposed. Conventional 2-D methods [13, 16], processing HSIs band by band, do not fully utilize the strong correlation between adjacent bands. 3-D methods such as Block-matching 4-D filtering (BM4D) [29], spectral-spatial adaptive hyperspectral total variation (SSAHTV) model [49], and sparse representation methods [44, 54] incorporate both spatial and spectral information and outperform conventional methods. However, those methods may need additional preprocessing to remove non-Gaussian noise.
@@
In real-world scenarios, HSIs are often contaminated by more than one type of noise due to atmospheric effects and instrument noise. Various methods have been proposed to remove the mixed noise, including low-rank matrix based methods, low-rank tensor based methods, and deep learning based methods. Low-rank matrix based methods [20, 37, 51] reshape an HSI into a matrix, e.g., each spectral band of the HSI is reshaped as each row of the matrix, and impose low-rankness on the reshaped HSI. Zhang et al. [52] formulated the HSI denoising problem as a low-rank matrix factorization problem and solved it by the “Go Decomposition” algorithm; Zhang et al. [51] proposed a double low-rank matrix decomposition which utilized the $\ell_{1}$ norm for the impulse noise and the matrix nuclear norm for stripes, and adopted augmented Lagrangian method (ALM) to solve the model; Yang et al. [47] also used double low-rankness, but added spatial-spectral total variation (SSTV) to the model and performed the decomposition on full band blocks (FBBs) rather than the entire HSI. However, imposing low-rankness on matrices can only exploit redundancy in two dimensions and may potentially lose important structural information due to reshaping.
@@
Low-rank tensor based methods [14, 18, 23, 33, 55] view HSIs as tensors, and can exploit redundancy in three or more dimensions. Tensor low-rank decompositions can preserve the spatial-spectral correlations and even nonlocal self-similarity if block matching is involved. Wang et al. [41] used Tucker tensor decomposition and an anisotropic SSTV regularization to characterize the piece-wise smooth structures of the HSI; Chen et al.  [12] proposed a low-rank tensor decomposition (LRTD) method, which utilized the higher-order singular value decomposition (HOSVD) for low-rankness and the matrix $\ell_{2,1}$ norm for characterizing the stripes, and adopted ALM for solving the optimization model. Cao et al. [6] employed the fact that spectral signatures of pixels lie in a low-dimensional subspace, and proposed a subspace-based nonlocal low-rank and sparse factorization (SNLRSF) method for removing mixed noise in HSI, which conducted nonlocal low-rank factorization via successive singular value decomposition (SVD). Xiong et al. [45] proposed the LRTFL0 method using low-rank block term decomposition and spectral-spatial $\ell_{0}$ gradient regularization to achieve gradient smoothness.
@@
For removing Gaussian noise, the existing low-rank based methods mainly utilize global spectral correlation of HSIs, and do not use its nonlocal self-similarity prior [5], except for SNLRSF. This limits the performance of restoring clean HSIs. Also, many of those methods use ALM to solve their low-rank regularization based models without providing any convergence analysis. In their ALM, the augmented Lagrangian subproblem is solved by alternating minimization via one Gauss–Seidel-type iteration, instead of an exact joint minimization. As a result, the convergence is not guaranteed for these nonconvex nonsmooth optimization problems with three or more blocks [10].
@@
Recently, some deep neural networks [4, 43, 50] have been proposed to denoise HSIs. Quasi-recurrent neural networks (QRNN) [4] combined recurrent neural networks (RNN) with convolutional neural networks (CNN); the 3-D version of QRNN (QRNN3D) [43], applying quasi-recurrent pooling function along spectrum, can effectively embed both structural spatial-spectral correlation and global correlation along spectrum of HSIs; CNN can also be used as a denoiser in a plug-and-play fashion for HSI denoising [38]. However, the performance of these methods may vary significantly depending on factors such as the datasets used, noise characteristics, and the size of the input data.
@@
(1) $\begin{split}\min_{\begin{subarray}{c}\mathcal{S},[\boldsymbol{X}_{1}],[\boldsymbol{X}_{2}],\\ [\boldsymbol{X}_{3}],[\mathcal{G}],\mathcal{L}\end{subarray}}\quad&\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\|\mathcal{S}\|^{p}_{2,p}+\|[\mathcal{G}]\|_{1,\boldsymbol{w}}\\ &+\frac{\delta}{2}\|\mathscr{R}(\mathcal{L})-[\mathcal{G}]\times_{1}[\boldsymbol{X}_{1}]\times_{2}[\boldsymbol{X}_{2}]\times_{3}[\boldsymbol{X}_{3}]\|^{2}_{F}\\ \mbox{s.t. }&[\boldsymbol{X}_{i}]^{\top}[\boldsymbol{X}_{i}]=[\boldsymbol{I}_{n_{i}}],\quad i=1,2,3,\end{split}$
@@
  • $\mathcal{S},\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $[\mathcal{G}]\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}\times N}$ and $[\boldsymbol{X}_{i}]\in\mathbb{R}^{m_{i}\times n_{i}\times N}$, $m_{i}\geq n_{i}$, $i=1,2,3$;
@@
    (2) $\|\mathcal{S}\|_{2,p}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\|\boldsymbol{s}_{:i_{2}i_{3}}\|^{p}_{2}\right)^{\frac{1}{p}}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\left(\sum_{i_{1}=1}^{I_{1}}{s}_{i_{1}i_{2}i_{3}}^{2}\right)^{\frac{p}{2}}\right)^{\frac{1}{p}},$
    and the tensor $\ell_{2,p}$ norm is exactly equal to the matrix $\ell_{2,p}$ norm of the unfolding matrix of ${\mathcal{S}}$ along the first dimension, i.e., $\|{\mathcal{S}}\|_{2,p}=\|\boldsymbol{S}_{(1)}\|_{2,p}$;
  • $\mathscr{R}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ denotes the block extraction operator that extracts blocks and forms a fourth order tensor, for example, the nonlocal similar FBB tensor, which can be regularized to exhibit low-rank properties;
  • $[\mathcal{G}]\times_{1}[\boldsymbol{X}_{1}]\times_{2}[\boldsymbol{X}_{2}]\times_{3}[\boldsymbol{X}_{3}]$ denotes an independent 3-D HOSVD (see the definition in section 2.2) with $[\mathcal{G}]$ being a stack of $N$ independent core tensors and $[\boldsymbol{X}_{i}]$ being a stack of $N$ independent mode-$i$ factor matrices, $i=1,2,3$, and $[\boldsymbol{X}_{i}]$ is independently orthogonal, i.e., $[\boldsymbol{X}_{i}]^{\top}[\boldsymbol{X}_{i}]=[\boldsymbol{I}_{n_{i}}]$, with $\boldsymbol{I}_{n_{i}}$ representing the identity matrix of size $n_{i}\times n_{i}$;
  • $\|[\mathcal{G}]\|_{1,\boldsymbol{w}}$ denotes the weighted tensor (component-wise) $\ell_{1}$ norm for a fourth order tensor $[\mathcal{G}]$ defined by
    (3) $\|[\mathcal{G}]\|_{1,\boldsymbol{w}}=\sum_{j=1}^{N}w_{j}\|[\mathcal{G}]^{(j)}\|_{1},$
    with $\boldsymbol{w}\in\mathbb{R}^{N}_{++}$ being a positive weight vector, and $[\mathcal{G}]^{(j)}$ being the $j$-th independent core tensor of $[\mathcal{G}]$.
Model (1) is a nonconvex nonsmooth minimization problem with orthogonal constraints. In particular, the first term of model (1) is a data fidelity term to remove Gaussian noise, the second term is a group sparsity measure to remove sparse noise with linear structures, and the last two terms are the sparsity-enhanced nonlocal low-rank tensor regularization terms.
@@
  • We propose a novel model to remove mixed noise in HSIs using a new sparsity-enhanced low-rank regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. For removing Gaussian noise, a tensor formed via the extraction operator $\mathscr{R}$ is regularized on its low-rankness using independent 3-D HOSVD with sparsity enhancement on its core tensors to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for removing the dead lines and stripes is extended from the matrix $\ell_{2,p}$ norm. Some new mathematical results on finding the proximal operator of the tensor $\ell_{2,p}$ norm are also presented. We show that model (1) has a nonempty and bounded solution set.
@@ 2.1 Notations
Throughout this paper, tensors are denoted by calligraphic uppercase letters, e.g., $\mathcal{X}$, matrices by boldface uppercase letters, e.g., $\boldsymbol{X}$, vectors by boldface lowercase letters, e.g., $\boldsymbol{x}$, scalars or entries by lowercase letters, e.g., $x$, and operators by curly uppercase letters, e.g., $\mathscr{R}$.
First, we introduce the notations for tensors. For a third order tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we let $x_{i_{1}i_{2}i_{3}}$ denote its $(i_{1},i_{2},i_{3})$-th entry, let $\boldsymbol{x}_{:i_{2}i_{3}}$ denote its $(i_{2},i_{3})$-th mode-$1$ fiber and let $\boldsymbol{X}_{::i_{3}}$ denote its $i_{3}$-th frontal slice. And $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$ means all its entries are positive. The mode-$k$ unfolding of a third order tensor $\mathcal{X}$ is denoted as $\boldsymbol{X}_{(k)}={\rm unfold}_{(k)}(\mathcal{X})$, which is the process to linearize all indexes except index $k$. The dimensions of $\boldsymbol{X}_{(k)}$ are $I_{k}\times\prod^{3}_{j=1,j\neq k}I_{j}$. An element $x_{i_{1}i_{2}i_{3}}$ of $\mathcal{X}$ corresponds to the position of $(i_{k},j)$ in matrix $\boldsymbol{X}_{(k)}$, where $j=1+\sum^{3}_{l=1,l\neq k}(i_{l}-1)\prod^{l-1}_{m=1,m\neq k}I_{m}$. The inverse process of the mode-$k$ unfolding of a tensor $\mathcal{X}$ is denoted by $\mathcal{X}={\rm fold}_{(k)}(\boldsymbol{X}_{(k)})$. An illustration of the unfoldings is presented in Fig. 2.
Figure 2: The unfoldings of tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$.
@@
Next, we provide the definitions for (limiting) subdifferentials and proximal operators. Let $f:\mathbb{R}^{d}\to(-\infty,+\infty]$ be a proper and lower semicontinuous function with a finite lower bound. The (limiting) subdifferential of $f$ at $\boldsymbol{x}\in\operatorname{dom}f:=\{\boldsymbol{x}\in\mathbb{R}^{d}|f(\boldsymbol{x})<\infty\}$, denoted by $\partial f(\boldsymbol{x})$, is defined as
$\partial f(\boldsymbol{x})\!:=\{\boldsymbol{u}\in\mathbb{R}^{d}|$ $\;\exists\boldsymbol{x}^{k}\to\boldsymbol{x},f(\boldsymbol{x}^{k})\to f(\boldsymbol{x})$
[…]
where ^$\hat{\partial}f(\boldsymbol{x})$ denotes the Fréchet subdifferential of $f$ at $\boldsymbol{x}\in\operatorname{dom}f$, which is the set of all $\boldsymbol{u}\in\mathbb{R}^{d}$ satisfying
(4) $\begin{split}\underset{\boldsymbol{y}\neq\boldsymbol{x},\boldsymbol{y}\rightarrow\boldsymbol{x}}{\lim\inf}\frac{f(\boldsymbol{y})-f(\boldsymbol{x})-\langle\boldsymbol{u},\boldsymbol{y}-\boldsymbol{x}\rangle}{\|\boldsymbol{y}-\boldsymbol{x}\|}\geq 0.\end{split}$
It follows from [32] that $\{\boldsymbol{u}\in\mathbb{R}^{d}|\exists\boldsymbol{x}^{k}\to\boldsymbol{x},f(\boldsymbol{x}^{k})\to f(\boldsymbol{x})\text{ and }\boldsymbol{u}^{k}\to\boldsymbol{u}\text{ with }\boldsymbol{u}^{k}\in\partial f(\boldsymbol{x}^{k})\mbox{ as }k\to\infty\}\subseteq\partial f(\boldsymbol{x})$. Next, the proximal operator of $f$ with parameter $\lambda>0$ evaluated at $\boldsymbol{x}\in\mathbb{R}^{d}$, denoted as $\text{prox}_{\lambda f}(\boldsymbol{x})$, is defined as
$\text{prox}_{\lambda f}(\boldsymbol{x}):=\underset{\boldsymbol{u}\in\mathbb{R}^{d}}{\arg\!\min}\left[f(\boldsymbol{u})+\frac{1}{2\lambda}\|\boldsymbol{u}-\boldsymbol{x}\|_{2}^{2}\right].$
Note that $\text{prox}_{\lambda f}$ is a single-valued map, if $f$ is a convex function. However, when $f$ is nonconvex, $\text{prox}_{\lambda f}(\boldsymbol{x})$ may have multiple points.
Also, we set $\mathbb{S}_{m,n}:=\{\boldsymbol{X}\in\mathbb{R}^{m\times n}|\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{n}\}$ as the Stiefel manifold with $m\geq n$ and set $\mathcal{T}_{\boldsymbol{X}}\mathbb{S}_{m,n}:=\{\boldsymbol{Y}\in\mathbb{R}^{m\times n}|\boldsymbol{Y}^{\top}\boldsymbol{X}+\boldsymbol{X}^{\top}\boldsymbol{Y}=\boldsymbol{0}\}$ as the tangent space of Stiefel manifold at $\boldsymbol{X}\in\mathbb{R}^{m\times n}$. We also set the Riemannian metric on Stiefel manifold as the metric induced from the Euclidean inner product. Then according to [1], the Riemannian gradient of a smooth function $f$ at $\boldsymbol{X}$ is given by
$\operatorname{grad}f(\boldsymbol{X}):=\operatorname{Proj}_{\mathcal{T}_{\boldsymbol{X}}\mathbb{S}_{m,n}}(\nabla f(\boldsymbol{X})),$
where $\operatorname{Proj}_{\mathcal{T}_{\boldsymbol{X}}\mathbb{S}_{m,n}}(\boldsymbol{Y}):=(\boldsymbol{I}_{m}-\boldsymbol{XX}^{\top})\boldsymbol{Y}+\frac{1}{2}\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{Y}-\boldsymbol{Y}^{\top}\boldsymbol{X})$.